### v1

@@ 3.1 Computational setup
Hard events are generated at leading order in QCD using a HPC-enabled version of the Sherpa event generator Gleisberg:2008ta , using the automated matrix-element generator COMIX Gleisberg_2008 and stored using the efficient and scalable HDF5 event-file format Hoche:2019flt ; Bothmann:2023ozs . [^1] Available at https://gitlab.com/hpcgen/me. For cross checks, we have also generated events with MadGraph5 Alwall_2014 using identical model parameters and phase-space cuts as the Sherpa samples, stored using the Les Houches Event File format Alwall:2006yp . All parton-multiplicity samples are generated with $10~{}\mathrm{M}$ events, except the highest multiplicity sample, which has $1~{}\mathrm{M}$ events of the process $e^{-}~{}p\rightarrow e^{-}~{}5j$. We consider a larger event sample unreasonable for the scope of this study, as the generation of unweighted events requires a significant investment of CPU resources (approximately $24~{}\mathrm{k}~{}\mathrm{CPUh}/1~{}\mathrm{M}~{}\text{events}$). With MadGraph5 we limit our study to events with up to $2$ additional partons, as we observed that the cross section became increasingly unstable with $3$ additional partons or more with the applied cuts. We have made all event samples publicly available Sample1 ; Sample2 ; Sample3 ; Sample4 ; Sample5 . The cuts are chosen very inclusively to cover multiple analyses and kinematic ranges with the cost of lower statistics in some multi-differential observables. The lowest-multiplicity sample is generated with a cut on virtuality $Q^{2}>2$ GeV2, while for higher multiplicities we further employ the $k_{\mathrm{T}}$-algorithm in the FastJet Cacciari_2012 implementation and apply an additional cut of $k_{\mathrm{T}}>2~{}\mathrm{GeV}$ to regulate infrared divergences.
@@
2. 2.)2 . ) $\mu_{\mathrm{R}}^{2}=\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle\right)\,.$ (22)
and their variations by a factor of $2$, which is a standard way of quantifying the uncertainty of theory calculations and Monte Carlo predictions. Here, $p_{\mathrm{T}}$ refers to the transverse momentum in the Breit frame and its average is taken as
@@
This scale choice takes into account events for which $Q^{2}$ is not necessarily the hardest energy scale of the process. When using a dynamic merging scale of eq. (19), our default choice of the reference scale is $\tilde{t}_{\mathrm{MS}}=(5~{}\mathrm{GeV})^{2}$ and $S=0.7$ for the $S$-parameter, which we have observed to yield good overall agreement with data. The dynamic merging scale is to be used with the standard choice of factorization scale $\mu_{\mathrm{F}}^{2}=Q^{2}$.
@@
  • CKKW-L with fixed merging scale and $\mu_{\mathrm{R}}^{2}=\mu_{\mathrm{F}}^{2}=\frac{1}{2}(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle)$ as in eq. (23),
  • UMEPS with fixed merging scale and $\mu_{\mathrm{R}}^{2}=\mu_{\mathrm{F}}^{2}=\frac{1}{2}(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle)$ as in eq. (23).
The scale choice of eq. (22) and the dynamical merging scale both contribute to the low-virtuality phase-space, and using both options led to too much of the desired effect. A setup with only one of these options lead to most realistic modelling in our studies, thus we only consider the three options listed above.
Parton-level events are merged with Pythia 8.3, using the Vincia antenna shower Brooks:2020upa in conjunction with its dedicated sectorized multi-jet merging framework Brooks:2020mab . The necessary modifications to enable merging in DIS will be included in a future Pythia release. The hadronization and fragmentation of showered configurations is performed with Pythia’s Lund string-fragmentation model Andersson:1983ia , using Vincia’s preliminary default tune Brooks:2020upa , which has been derived based on the methods developed for Pythia’s Monash 2013 tune Skands:2014pea . In the shower evolution, we use a two-loop running of the strong coupling with $\alpha_{\mathrm{S}}(m_{Z})=0.118$ in the $\overline{\mathrm{MS}}$ scheme. The effective coupling is evaluated in the CMW scheme Catani:1990rr with additional renormalisation-scale factors Brooks_2020 . Analyses are performed using Pythia’s internal analysis routines and the Rivet framework Bierlich:2019rhm , for which we have prepared new analyses for HERA data. The details of these analyses are discussed in section 4.
@@ 3.2 Validation
Figure 4: Merged cross sections $\sigma^{\mathrm{merged}}_{4}$ of CKKW-L (red) and UMEPS (green) algorithms in $+4$-jet merging as ratios to the Born level cross section $\sigma_{0}$. Both algorithms use the fixed merging scale and $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle\right)$. Cross section ratios are presented as a function of merging scale $\tilde{t}_{\mathrm{MS}}$.
The exact definition of a merging scale should not affect the full inclusive merged cross section too much, as it dictates the ratios of different contributions of the full cross section. Just as is the case for factorization and renormalization scales, the final cross sections should ideally not be too sensitive to different scale choices. New parameters are introduced in merging algorithms, and the effects of varying these parameters should also be a part of the uncertainty analysis in multi-jet merging predictions. The effect of varying the fixed value of $\tilde{t}_{\mathrm{MS}}$ is shown in figure 5 for the three considered setups. The $2\rightarrow 1$ resolution scale (or differential jet rate), $k_{\mathrm{T},2\rightarrow 1}$, is the transverse momentum scale at which the event transitions from a two-jet state to a one-jet state. It is related to the value of a parton shower evolution parameter at a 1-to-2 splitting. The full results from each setup (black lines) are the sums of contributions from each event sample, with the $+0$-jet $(e^{-}~{}p\rightarrow e^{-}~{}j)$ sample in red, $+1$-jet $(e^{-}~{}p\rightarrow e^{-}~{}2j)$ in blue and $+2$-jet $(e^{-}~{}p\rightarrow e^{-}~{}3j)$ in orange. Variations of the merging scale parameter are plotted as dashed lines for $(9~{}\mathrm{GeV})^{2}$, dotted lines for $(3~{}\mathrm{GeV})^{2}$ and the central value $(5~{}\mathrm{GeV})^{2}$ as solid lines. Figure 5 shows how similar results are obtained with different setups, and how the value of the merging scale affects each setup. We find that while contributions from each multiplicity are sensitive to the $\tilde{t}_{\mathrm{MS}}$ choice, especially in case of a fixed merging scale, the total cross section is always less affected. The contributions $\widehat{B}_{n}-\int\widehat{B}_{n+1\rightarrow n}-...-\int\widehat{B}_{n+1\rightarrow 0}$ for UMEPS merging can be compared to contributions $B_{n}$ of CKKW-L merging. With a dynamic merging scale, small-$Q^{2}$ dynamics are less sensitive to $\tilde{t}_{\mathrm{MS}}$, while the fixed merging results vary considerably more. This can be explained with the weak sensitivity of $t_{\mathrm{MS}}$ to $\tilde{t}_{\mathrm{MS}}$ at low-$Q^{2}$ region as shown in figure 3. Using a fixed merging scale results to a high variation across jet rates in lower values of $k_{\mathrm{T}}$. This effect holds even when restoring unitarity with the UMEPS algorithm. However, the experimental data confronted in section 4 have cuts on jet transverse momenta $p_{\mathrm{T}}\sim 4$ GeV, which roughly translates to $\log_{10}(k_{\mathrm{T}}/\mathrm{GeV})\sim 0.6$, so the high variation in low-$k_{\mathrm{T}}$ region is not present in the final results. Significantly reduced variations of the individual contributions in case of dynamical merging-scale determination makes it clearly an appropriate choice for multi-jet merging in DIS. The combined jet rates are, however, still fairly similar in all of the studied cases and thus also the algorithms with fixed merging scale can be considered as applicable options.
@@
Figure 5: Jet rates with variation in $\tilde{t}_{\mathrm{MS}}$, using CKKW-L with a dynamic merging scale and $\mu_{\mathrm{F}}^{2}=Q^{2}$ (left) and CKKW-L (middle) and UMEPS with a fixed merging scale and $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle\right)$ (right) with $+2$ jets. The full merged results (black line) are the sums of contributions from the jet multiplicity samples (colored lines), with variations in the parameter $\tilde{t}_{\mathrm{MS}}$ shown as dashed and dotted lines.
@@ 4.1 Jet multiplicities
Measurements of multi-jet cross sections have also been carried out in the analysis. Dijet distributions of figure 7 are calculated with the average $p_{\mathrm{T}}$ of the hardest two jets. Merging of two additional jets is again sufficient, with additional samples contributing only little to the cross section. The merging setups with fixed $\tilde{t}_{\mathrm{MS}}$ fail to describe the high-$p_{\mathrm{T}}$ tail of the data, especially with the UMEPS algorithm, while CKKW-L with dynamic merging scale provides a good description of the data also in this region, though some statistical fluctuations are still present.
@@
Figure 6: Jet production cross sections in terms of inclusive jet transverse momenta in the range $3022<Q^{2}<$ 30 GeV2. Different amount of additional final state partons are shown in the three merging setups: dynamic CKKW-L merging with $\mu_{\mathrm{F}}^{2}=Q^{2}$ and fixed CKKW-L and UMEPS merging with $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle\right)$. Data from H1 H1:2016goa .
@@ 4.2 Merging scale variations
For a sufficient description of jets at very low virtualities, a dynamic merging scale allows the matrix element contributions to access a larger fraction of phase space than in case of a fixed merging scale. The effect of varying the merging scale is presented in figures 9-11 for each of the three setups for three different $Q^{2}$ bins ranging from $5.5~{}\mathrm{GeV}^{2}$ to $80~{}\mathrm{GeV}^{2}$ for inclusive jets, dijets and trijets. The resulting cross sections are somewhat sensitive to parameter values related to the merging scale. In case of dynamic merging scale, panels a), we vary $S$ and $\tilde{t}_{\mathrm{MS}}$ of eq. (19) within ranges of $S=0.7\pm 0.1$ and $\tilde{t}_{\mathrm{MS}}=(5^{+4}_{-2}~{}\mathrm{GeV})^{2}$. Variation of the $S$ parameter has a 20% effect in the inclusive jet cross section but is reduced to less than 10% in higher virtuality events where the value of $t_{\mathrm{MS}}$ is close to $\tilde{t}_{\mathrm{MS}}$, as evident from figure 3. Variation of the $\tilde{t}_{\mathrm{MS}}$ parameter, in turn, has smaller impact at low-$Q^{2}$ bins but becomes dominant at larger values of $Q^{2}$. On average, the combined effect from these variations are around $40$% for inclusive jet cross section and dijets and slightly smaller for trijets.
@@
Figure 9: Inclusive jet cross sections with 4 additional jets in a) dynamic CKKW-L merging with $\mu_{\mathrm{F}}^{2}=Q^{2}$, b) fixed CKKW-L merging with $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle\right)$ and c) fixed UMEPS merging with $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+\langle p_{\mathrm{T}}^{2}\rangle\right)$. Light-red error bands show the variation in $\tilde{t}_{MS}=(5^{+4}_{-3}~{}\mathrm{GeV})^{2}$, and the diagonal line pattern shows the variation in parameter $S=0.7\pm 0.1$. Statistical uncertainties are shown as a vertical bar in each bin. Data and MC results have been scaled by a factor of $100$ for $85.5<Q^{2}<$ 8 GeV2 and $10$ for $3022<Q^{2}<$ 30 GeV2. Data from H1 H1:2016goa .
@@ 4.3 Factorization and renormalization scale variations
Figure 12: Comparison of $+2$-jet CKKW-L dynamic merging on parton level events from Sherpa and MadGraph5. Jet cross sections in terms of jet transverse energy by photon virtuality in central pseudorapidity region (left), and in terms of only transverse energy in low-$Q^{2}$ region (right). Uncertainty bands from factorization and renormalization scale variations. Data from H1 H1:2002qhb .

### v3

@@ 3.1 Computational setup
Hard events are generated at leading order in QCD using a HPC-enabled version of the Sherpa event generator Gleisberg:2008ta , using the automated matrix-element generator COMIX Gleisberg_2008 and stored using the efficient and scalable HDF5 event-file format Hoche:2019flt ; Bothmann:2023ozs . [^1] Available at https://gitlab.com/hpcgen/me. For cross checks, we have also generated events with MadGraph5 Alwall_2014 using identical model parameters and phase-space cuts as the Sherpa samples, stored using the Les Houches Event File format Alwall:2006yp . All parton-multiplicity samples are generated with $10~{}\mathrm{M}$ events, except the highest multiplicity sample, which has $1~{}\mathrm{M}$ events of the process $e^{-}~{}p\rightarrow e^{-}~{}5j$. We consider a larger event sample unreasonable for the scope of this study, as the generation of unweighted events requires a significant investment of CPU resources (approximately $24~{}\mathrm{k}~{}\mathrm{CPUh}/1~{}\mathrm{M}~{}\text{events}$). With MadGraph5 we limit our study to events with up to $2$ additional partons, as we observed that the cross section became increasingly unstable with $3$ additional partons or more with the applied cuts. We have made all event samples publicly available Sample1 ; Sample2 ; Sample3 ; Sample4 ; Sample5 . The cuts are chosen very inclusively to cover multiple analyses and kinematic ranges with the cost of lower statistics in some multi-differential observables. The lowest-multiplicity sample is generated with a cut on virtuality $Q^{2}>2$ GeV2, while for higher multiplicities we further employ the longitudinally invariant inclusive $k_{\mathrm{T}}$ algorithm in the FastJet Cacciari_2012 implementation, in both the analyses and event generators, and apply an additional cut of $k_{\mathrm{T}}>2~{}\mathrm{GeV}$ in the center-of-momentum frame of the beams to regulate infrared divergences.
@@
2. 2.)2 . ) $\mu_{\mathrm{R}}^{2}=\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+H_{\mathrm{T}}^{2}\right)\,.$ (29)
and their variations by a factor of $2$, which is a standard way of quantifying the uncertainty of theory calculations and Monte Carlo predictions. Here, $p_{\mathrm{T}}$ refers to the transverse momentum in the Breit frame and its scalar sum is taken as
@@
This scale choice takes into account events for which $Q^{2}$ is not necessarily the hardest energy scale of the process. Notice that in case of only one outgoing parton $H_{\mathrm{T}}=0$ in the Breit frame. When using a dynamic merging scale of eq. (26), our default choice of the reference scale is $\tilde{t}_{\mathrm{MS}}=(5~{}\mathrm{GeV})^{2}$ and $S=0.7$ for the $S$-parameter, which we have observed to yield good overall agreement with data. The dynamic merging scale is to be used with the standard choice of factorization scale $\mu_{\mathrm{F}}^{2}=Q^{2}$.
@@
  • CKKW-L with fixed merging scale and $\mu_{\mathrm{R}}^{2}=\mu_{\mathrm{F}}^{2}=\frac{1}{2}(Q^{2}+H_{\mathrm{T}}^{2})$ as in eq. (29),
  • UMEPS with fixed merging scale and $\mu_{\mathrm{R}}^{2}=\mu_{\mathrm{F}}^{2}=\frac{1}{2}(Q^{2}+H_{\mathrm{T}}^{2})$ as in eq. (29).
The scale choice of eq. (29) and the dynamical merging scale both contribute to the low-virtuality phase space, and using both options led to too much of the desired effect. A setup with only one of these options lead to most realistic modelling in our studies, thus we only consider the three options listed above.
Parton-level events are merged with Pythia 8.3, using the Vincia antenna shower Brooks:2020upa in conjunction with its dedicated sectorized multi-jet merging framework Brooks:2020mab . The necessary modifications to enable merging in DIS are included in Pythia version 8.313 onwards. The hadronization and fragmentation of showered configurations is performed with Pythia’s Lund string-fragmentation model Andersson:1983ia , using Vincia’s preliminary default tune Brooks:2020upa , which has been derived based on the methods developed for Pythia’s Monash 2013 tune Skands:2014pea . In the shower evolution, we use a two-loop running of the strong coupling with $\alpha_{\mathrm{S}}(m_{Z})=0.118$ in the $\overline{\mathrm{MS}}$ scheme. The effective coupling is evaluated in the CMW scheme Catani:1990rr with additional renormalization-scale factors Brooks_2020 . Analyses are performed using Pythia’s internal analysis routines and the Rivet framework Bierlich:2019rhm , for which we have prepared new analyses for HERA data. The details of these analyses are discussed in section 4.
Figure 3: Merged cross sections $\sigma^{\mathrm{merged}}_{4}$ of CKKW-L (red) and UMEPS (green) algorithms in $+4$-jet merging as ratios to the Born level cross section $\sigma_{0}$. Both algorithms use the fixed merging scale and $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+H_{\mathrm{T}}^{2}\right)$. Cross section ratios are presented as a function of merging scale $\tilde{t}_{\mathrm{MS}}$.
@@ 3.2 Validation
Figure 4: Jet rates with variation in $\tilde{t}_{\mathrm{MS}}$, using CKKW-L with a dynamic merging scale and $\mu_{\mathrm{F}}^{2}=Q^{2}$ (left) and CKKW-L (middle) and UMEPS with a fixed merging scale and $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+H_{\mathrm{T}}^{2}\right)$ (right) with $+2$ jets. The full merged results (black line) are the sums of contributions from the jet multiplicity samples (colored lines), with variations in the parameter $\tilde{t}_{\mathrm{MS}}$ shown as dashed and dotted lines.
The exact definition of a merging scale should not affect the full inclusive merged cross section too much, as it dictates the ratios of different contributions of the full cross section. Just as is the case for factorization and renormalization scales, the final cross sections should ideally not be too sensitive to different scale choices. New parameters are introduced in merging algorithms, and the effects of varying these parameters should also be a part of the uncertainty analysis in multi-jet merging predictions. The effect of varying the fixed value of $\tilde{t}_{\mathrm{MS}}$ is shown in figure 4 for the three considered setups. The $2\rightarrow 1$ resolution scale (or differential jet rate), $k_{\mathrm{T},2\rightarrow 1}$, is the transverse momentum scale at which the event transitions from a two-jet state to a one-jet state. It is related to the value of a parton-shower evolution parameter at a 1-to-2 splitting. The full results from each setup (black lines) are the sums of contributions from each event sample, with the $+0$-jet $(e^{-}~{}p\rightarrow e^{-}~{}j)$ sample in red, $+1$-jet $(e^{-}~{}p\rightarrow e^{-}~{}2j)$ in blue and $+2$-jet $(e^{-}~{}p\rightarrow e^{-}~{}3j)$ in orange. Variations of the merging-scale parameter are plotted as dashed lines for $(9~{}\mathrm{GeV})^{2}$, dotted lines for $(3~{}\mathrm{GeV})^{2}$ and the central value $(5~{}\mathrm{GeV})^{2}$ as solid lines. Figure 4 shows how similar results are obtained with different setups, and how the value of the merging scale affects each setup. We find that while contributions from each multiplicity are sensitive to the $\tilde{t}_{\mathrm{MS}}$ choice, especially in case of a fixed merging scale, the total cross section is always less affected. The contributions $\widehat{\mathrm{B}}_{n}-\int\widehat{\mathrm{B}}_{n+1\rightarrow n}-...-\int\widehat{\mathrm{B}}_{n+1\rightarrow 0}$ for UMEPS merging can be compared to contributions $\mathrm{B}_{n}$ of CKKW-L merging. With a dynamic merging scale, small-$Q^{2}$ dynamics are less sensitive to $\tilde{t}_{\mathrm{MS}}$, while the fixed merging results vary considerably more. This can be explained with the weak sensitivity of $t_{\mathrm{MS}}$ to $\tilde{t}_{\mathrm{MS}}$ at low-$Q^{2}$ region as shown in figure 5. Using a fixed merging scale results to a high variation across jet rates in lower values of $k_{\mathrm{T}}$. This effect holds even when restoring unitarity with the UMEPS algorithm. However, the experimental data confronted in section 4 have cuts on jet transverse momenta $p_{\mathrm{T}}\sim 4$ GeV, which roughly translates to $\log_{10}(k_{\mathrm{T}}/\mathrm{GeV})\sim 0.6$, so the high variation in low-$k_{\mathrm{T}}$ region is not present in the final results. Significantly reduced variations of the individual contributions in case of dynamical merging-scale determination makes it clearly an appropriate choice for multi-jet merging in DIS. The combined jet rates are, however, still fairly similar in all of the studied cases and thus also the algorithms with fixed merging scale can be considered as applicable options.
@@ 4 Results
Figure 6: Jet production cross sections in terms of inclusive jet transverse momenta in the range $3022<Q^{2}<$ 30 GeV2. Different amount of additional final state partons are shown in the three merging setups: dynamic CKKW-L merging with $\mu_{\mathrm{F}}^{2}=Q^{2}$ and fixed CKKW-L and UMEPS merging with $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+H_{\mathrm{T}}^{2}\right)$. Data from H1 H1:2016goa .
@@ 4.1 Jet multiplicities
Measurements of multi-jet cross sections have also been carried out in the analysis. Dijet distributions of figure 7 are calculated with the average $p_{\mathrm{T}}$ of the hardest two jets. Merging of two additional jets is again sufficient, with additional samples contributing only little to the cross section. The merging setups with fixed $\tilde{t}_{\mathrm{MS}}$ fail to describe the $p_{\mathrm{T}}$-slope of the data, overshooting it in the low-$p_{\mathrm{T}}$ region and undershooting in the high-$p_{\mathrm{T}}$ region. The CKKW-L merging with dynamic merging scale, instead, provides a good description of the data throughout the considered $p_{\mathrm{T}}$ range, though some statistical fluctuations are still present.
@@ 4.2 Merging scale variations
For a sufficient description of jets at very low virtualities, a dynamic merging scale allows the matrix element contributions to access a larger fraction of phase space than in case of a fixed merging scale. The effect of varying the merging scale is presented in figures 9-11 for each of the three setups for three different $Q^{2}$ bins ranging from $5.5~{}\mathrm{GeV}^{2}$ to $80~{}\mathrm{GeV}^{2}$ for inclusive jets, dijets and trijets. The resulting cross sections are somewhat sensitive to parameter values related to the merging scale. In case of dynamic merging scale, panels a), we vary $S$ and $\tilde{t}_{\mathrm{MS}}$ of eq. (26) within ranges of $S=0.7\pm 0.1$ and $\tilde{t}_{\mathrm{MS}}=(5^{+4}_{-2}~{}\mathrm{GeV})^{2}$. Variation of the $S$-parameter has a 20% effect in the inclusive jet cross section but is reduced to less than 10% in higher virtuality events where the value of $t_{\mathrm{MS}}$ is close to $\tilde{t}_{\mathrm{MS}}$, as evident from figure 5. Variation of the $\tilde{t}_{\mathrm{MS}}$-parameter, in turn, has smaller impact at low-$Q^{2}$ bins but becomes dominant at larger values of $Q^{2}$. On average, the combined effect from these variations are around $40$% for inclusive jet cross section and dijets and slightly smaller for trijets.
@@
Figure 9: Inclusive jet cross sections with 4 additional jets in a) dynamic CKKW-L merging with $\mu_{\mathrm{F}}^{2}=Q^{2}$, b) fixed CKKW-L merging with $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+H_{\mathrm{T}}^{2}\right)$ and c) fixed UMEPS merging with $\mu_{\mathrm{F}}^{2}=\frac{1}{2}\left(Q^{2}+H_{\mathrm{T}}^{2}\right)$. Light-red error bands show the variation in $\tilde{t}_{MS}=(5^{+4}_{-2}~{}\mathrm{GeV})^{2}$, and the diagonal line pattern shows the variation in parameter $S=0.7\pm 0.1$. Statistical uncertainties are shown as a vertical bar in each bin. Data and MC results have been scaled by a factor of $100$ for $85.5<Q^{2}<$ 8 GeV2 and $10$ for $3022<Q^{2}<$ 30 GeV2. Data from H1 H1:2016goa .
@@ 4.3 Factorization and renormalization scale variations
Figure 12: Comparison of $+2$-jet CKKW-L dynamic merging on parton level events from Sherpa and MadGraph5. Jet cross sections in terms of jet transverse energy by photon virtuality in different pseudorapidity regions. Uncertainty bands from factorization and renormalization scale variations. Data from H1 H1:2002qhb .